OpenStudy (anonymous):

Rewrite the expression 1/2ln(x – 6) + 3ln(x +1) as a single logarithm.

OpenStudy (anonymous):

\[\ln (x-6)^{1/2}+\ln (x+1)^{3}\]

OpenStudy (anonymous):

\[\ln \sqrt{x-6}*(x+1)^{3}\]

OpenStudy (anonymous):